# Hybrid-pair superfluidity in a strongly driven Fermi gas

### v1

@@ Abstract
We explore the paired superfluid phases of a Fermi gas in the presence of a continuous Rabi drive. We focus on the case where two components are strongly coupled by the drive, forming hybrid superpositions, and interacting with an uncoupled third component. Using a generalized Bardeen-Cooper-Schrieffer (BCS) ansatz, we show that there are two coupled superfluid order parameters, and we obtain the associated free energy and quasiparticle excitation spectrum. We find that we can drive BCS-BCS, BCS-Bose-Einstein condensate (BEC) and BEC-BEC crossovers purely by varying the detuning of the Rabi drive from the bare transition, with the precise crossover depending on the sign of the underlying interactions between the coupled and uncoupled components. We furthermore identify an exotic excited branch which features both normal to BCS superfluid transitions, as well as a BCS-BEC-BCS crossover. Introducing a generalized Thouless criterion, we show that this behavior is reflected in the critical temperature for superfluidity. Our Rabi-coupled scenario also gives possesses additional thermodynamic properties related to the pseudospin of the coupled components, which provide novel signatures of the state of the many-body system. The Rabi-driven Fermi gas thus emerges as a unique platform for engineering and probing a rich array of multi-band superfluid phases.
@@ I Introduction
Motivated by these developments, here we investigate the superfluid properties of a strongly interacting, strongly Rabi-coupled Fermi gas. The central idea is illustrated in Fig. 1. A strong rf field creates a superposition (purple) of two atomic states ($|1\rangle$=red and $|2\rangle$=blue), where each state interacts differently with a third atomic species ($|3\rangle$=gray). Taking the total number of atoms in the Rabi-driven components equal to the third component, i.e., $N_{1}+N_{2}=N_{3}$, we find that there is the prospect of driving a rich array of different superfluid crossovers—BCS-BCS, BCS-BEC, and BEC-BEC—purely by varying the detuning of the rf field from the bare transition. While our proposed setup naturally involves three states, we emphasize that for sufficiently large Rabi coupling it effectively becomes a two-component Fermi gas where the upper Rabi-dressed state is minimally occupied. In this sense, our focus is distinct from previous theory works on three-component Fermi gases, such as those exploring analogs of color superfluidity [33, 34, 35, 36, 37, 38, 39, 40].
@@
Figure 1: (a) Rabi-dressed atom (purple) that is obtained by driving the transition between state |$\ket{1}$ (red) and state |$\ket{2}$ (blue). (b) Both states |$\ket{1}$ and |$\ket{2}$ can interact with a third component (gray) via short range interactions that are characterized by the scattering lengths $a_{13}$ and $a_{23}$, respectively. (c) Illustration of the four different crossovers that can be accessed by the ground state of the Rabi driven Fermi gas when varying the detuning $\delta$, depending on the signs of the scattering lengths. The hues of the Rabi-dressed atoms indicate their spin composition.
@@ II Model
Here, $\hat{f}_{{\bf k}j}$ ($\hat{f}_{{\bf k}j}^{\dagger}$) are annihilation (creation) operators for fermions of species $j=\{1,2,3\}$. The states $|1\rangle$ and $|2\rangle$ are coupled via a Rabi drive of strength $\Omega$ which is detuned by $\delta$ with respect to the bare $|1\rangle$-$|2\rangle$ transition. We therefore take the dispersions for each species at momentum ${\bf k}$ to be $\epsilon_{{\bf k}1}\equiv\epsilon_{{\bf k}}=k^{2}/2m$, $\epsilon_{{\bf k}2}=\epsilon_{\bf k}+\delta$ and $\epsilon_{{\bf k}3}=k^{2}/2m_{3}$. We take a general mass ratio $m_{3}/m$ for species $|3\rangle$ compared with the coupled species $|1\rangle$ and $|2\rangle$, since many of our results for the paired superfluid phases only depend on the reduced mass $m_{r}=mm_{3}/(m+m_{3})$. They thus apply to general mass-imbalanced Fermi mixtures [42], as well as equal-mass 6Li or 40K gases where the three species would represent three distinct hyperfine states. For instance, the scenario where three hyperfine components of 6Li are simultaneously present has already been experimentally realized in Refs. [43, 44, 45, 46, 47]. Throughout, we work in units where the reduced Planck constant $\hbar$, the Boltzmann constant $k_{\rm B}$, and the volume are all set to 1.
In this work, we only consider $|1\rangle$-$|3\rangle$ and $|2\rangle$-$|3\rangle$ interactions, while we assume that the $|1\rangle$-$|2\rangle$ interactions are negligible, a scenario which can be achieved in 6Li [48]. To model the interactions, we use a short-range contact potential leading to Eq. (2) in momentum space. The bare interaction strengths $g_{j3}$ between species |$\ket{j}$ and $|3\rangle$ are related to the scattering lengths $a_{j3}$ via the renormalization condition
@@ III BCS mean field theory
$[\hat{H},\hat{N}_{3}]=0,\leavevmode\nobreak\ \leavevmode\nobreak\ [\hat{H},\hat{N}_{1}+\hat{N}_{2}]=0.$ (9)
@@ III.1 Generalized BCS ansatz
|$\ket{\psi}$ $=\prod_{\bf k}(u_{\bf k}+v_{{\bf k}}\hat{f}_{{\bf k}3}^{\dagger}\hat{f}_{-{\bf k}a}^{\dagger})\ket{0},$ (10)
where $u_{\bf k}$ and $v_{\bf k}$ are variational coefficients satisfying $|u_{\bf k}|^{2}+|v_{\bf k}|^{2}=1$ as in the standard BCS theory. In writing this ansatz, we have introduced the Fermi operator $\hat{f}_{{\bf k}a}$ which is related to the operators of species |$\ket{1}$ and |$\ket{2}$ via the transformation
$\begin{pmatrix}\hat{f}_{{\bf k}a}\\ \hat{f}_{{\bf k}s}\end{pmatrix}$ $=\begin{pmatrix}c_{\bf k}&&-s_{\bf k}\\ s_{\bf k}&&c_{\bf k}\end{pmatrix}\begin{pmatrix}\hat{f}_{{\bf k}1}\\ \hat{f}_{{\bf k}2}\end{pmatrix},$ (11)
with $|c_{\bf k}|^{2}+|s_{\bf k}|^{2}=1$. Note that $c_{\bf k}$, $s_{\bf k}$ are variational coefficients that define the distribution between the two pairing states and thus remain to be determined [^1] In general, once the system is interacting we have $\hat{f}_{{\bf k}a/s}\neq\hat{f}_{{\bf k}\mp}$..
@@ III.2 Gap equations
00 0 $=\zeta_{\bf k}u_{\bf k}v_{\bf k}+(u_{\bf k}^{2}-v_{\bf k}^{2})\left(\Delta_{1}c_{\bf k}-\Delta_{2}s_{\bf k}\right),$ (14a)
00 0 $=\left[2\delta c_{\bf k}s_{\bf k}-\Omega(c_{\bf k}^{2}-s_{\bf k}^{2})\right]v_{\bf k}^{2}-2u_{\bf k}v_{\bf k}\left(\Delta_{1}s_{\bf k}+\Delta_{2}c_{\bf k}\right).$ (14b)
Here, we have introduced two order parameters $\Delta_{1}$ and $\Delta_{2}$ defined by
@@
Following the usual BCS theory and solving Eq. (14a) for $\theta_{\bf k}$, we obtain $\tan{2\theta_{\bf k}}=-2(\Delta_{1}c_{\bf k}-\Delta_{2}s_{\bf k})/\zeta_{\bf k},$ from which we can deduce
$v_{\bf k}^{2}=\frac{1}{2}\left(1-\frac{\zeta_{\bf k}}{2E_{\bf k}}\right),\leavevmode\nobreak\ \leavevmode\nobreak\ u_{\bf k}v_{\bf k}=-\frac{\Delta_{1}c_{\bf k}-\Delta_{2}s_{\bf k}}{2E_{\bf k}},$ (16)
@@
×$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \times\frac{\left(t_{\bf k}^{2}+2t_{\bf k}\frac{\delta}{\Omega}-1\right)}{4\left(t_{\bf k}+\frac{\Delta_{2}}{\Delta_{1}}\right)^{2}\left(1-t_{\bf k}\frac{\Delta_{2}}{\Delta_{1}}\right)}=\frac{\Delta_{1}^{2}}{\Omega^{2}},$ (22)
@@
In our model, the solution with $\text{sgn}(\Delta_{2}/\Delta_{1})=-1$ corresponds to the stable ground state, while the one with $\text{sgn}(\Delta_{2}/\Delta_{1})=+1$ corresponds to an excited state, which appears as a saddle point of the free energy. This is illustrated in Fig. 2, which shows colormaps of $\mathcal{F_{\text{BCS}}}(\Delta_{1},\Delta_{2})$. Indeed, we can observe in panel (a) the existence of 4 stationary points corresponding to a pair of minima with $\text{sgn}(\Delta_{2}/\Delta_{1})=-1$, and a pair of saddle points appearing in the orthogonal direction in the $(\Delta_{1},\Delta_{2})$ plane. Furthermore, we see that the excited branch is topologically distinct from the ground state, in the sense that one of the order parameters must go through zero when evolving from one state to another. This suggests the possibility of tuning into a many-body excited state that is long lived compared to the typical timescales of the system.
@@
We also observe that the two types of BCS solutions are connected to the two-body bound states in the limit of vanishing density (see Appendix A). However, we stress that while it has been shown that the existence of two bound states requires both scattering lengths to be positive [41], we find that this is not a necessary condition for two solutions of the order parameters to exist in the many-body regime. In Section IV, we discuss how one can use a generalized Thouless criterion to investigate the existence of the metastable excited state in the many-body regime.
@@ III.3 Quasiparticle excitations
which have the property $\hat{\gamma}_{{\bf k}\sigma}\ket{\psi}=0$, where $\sigma\in\{\uparrow,\downarrow\}$, so that the BCS state acts as the vacuum of Bogoliubov excitations. To determine the quasiparticle energies we first calculate the following expectation values [57]
@@ IV Thouless criterion
$\leavevmode\nobreak\ \leavevmode\nobreak\ -\left(\Delta_{1}\quad\Delta_{2}\right)\mathbf{T}^{\text{mb}}(0)^{-1}\begin{pmatrix}\Delta_{1}\\ \Delta_{2}\end{pmatrix}+\mathcal{O}(\Delta_{i}^{2}\Delta_{j}^{2}).$ (36)
@@
This corresponds to a generalized version of the Thouless criterion [58]. Solving Eq. (38) together with the number equations in Eq. (35) allows us to determine the critical temperature at fixed density. Specifically, renormalizing the diagonal elements of the inverse many-body $T$ matrix in Eq. (37) in the same manner as the gap equation in Eq. (19), we find that the Thouless criterion takes the form
@@
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ +\left(s^{2}\frac{m_{r}}{2\pi a_{13}}+c^{2}\frac{m_{r}}{2\pi a_{23}}\right)I_{-}+I_{+}I_{-}=0,$ (39)
@@ IV.1 Thouless criterion in the limit of zero temperature
We now wish to evaluate the generalized Thouless criterion in the limit of zero temperature and weak interactions. To evaluate $I_{\pm}$ analytically in this limit, we assume equal masses $m=m_{3}$ and a large Rabi drive $\Omega$ compared with the Fermi energy $E_{F}$. This implies that for weak interactions we have $\mu-\epsilon_{-}\simeq\mu_{3}\simeq E_{F}$. In this case, $I_{-}$ becomes independent of the parameters of the Rabi drive and takes the form
$I_{-}$ $=\sum_{{\bf k}}\left[\frac{\tanh(\beta(\epsilon_{\bf k}-E_{F})/2)}{2(\epsilon_{\bf k}-E_{F})}-\frac{1}{2\epsilon_{\bf k}}\right]$
$\simeq\frac{mk_{F}}{2\pi^{2}}\ln\left(\frac{8e^{\gamma-2}\beta E_{F}}{\pi}\right).$ (41)
Here, $\gamma\simeq 0.577$ is the Euler constant, and to obtain the second line we have used the fact that $I_{-}$ diverges logarithmically as temperature is taken to zero, thus allowing us to analytically capture the leading divergence [59]. By contrast, $I_{+}$ is finite as $T\to 0$ and depends on the Rabi-drive parameters:
$I_{+}$ $\simeq\sum_{{\bf k}}\left[\frac{1-\Theta(E_{F}-\epsilon_{\bf k})}{2\epsilon_{\bf k}-2E_{F}+\epsilon_{+}-\epsilon_{-}}-\frac{1}{2\epsilon_{\bf k}}\right]$ (42)
$=-\frac{mk_{F}}{2\pi^{2}}\left[1+\sqrt{\frac{\epsilon_{+}-\epsilon_{-}}{2E_{F}}-1}\arctan\sqrt{\frac{\epsilon_{+}-\epsilon_{-}}{2E_{F}}-1}\right],$
$s^{2}\frac{m}{4\pi a_{13}}+c^{2}\frac{m}{4\pi a_{23}}+I_{+}=0.$ (43)
Since $I_{+}<0$ is an even and concave function of $\delta$, the potential solutions of Eq. (51) will appear pairwise if $s^{2}/a_{13}+c^{2}/a_{23}\geq 4\pi|I_{+}|/m$ is realized at some detunings. Thus, such solutions can only occur when at least one of the scattering lengths is positive. A pair of $T=0$ solutions can be clearly seen in Fig. 3(a), at the points where the excited state $T^{*}$ goes to zero. The paired solutions of Eq. (51) thus define the critical detunings in-between which the excited state exists for a given set of scattering lengths.
IV.2 Thouless criterion at finite temperature
$T^{*}/T_{F}=\frac{8e^{\gamma-2}}{\pi}e^{F(\delta,\Omega,E_{F},a_{13},a_{23})}.$ (44)
$F=\frac{\pi}{2k_{F}}\frac{1+\frac{4\pi}{m}I_{+}(a_{13}s^{2}+a_{23}c^{2})}{a_{13}c^{2}+a_{23}s^{2}+\frac{4\pi}{m}I_{+}a_{13}a_{23}},$ (45)
where, for simplicity, we again take $m_{3}=m$. In the limit $E_{F}/\Omega\ll 1$, this function reduces to $F\simeq\frac{\pi}{2k_{F}a_{-}}$, with
the two-body scattering length between a particle of species $|3\rangle$ and the lower dressed particle $\hat{f}_{{\bf k}-}$ [see Eq. (5)], as introduced in Ref. [41]. This result for the critical temperature closely matches the standard result for the critical temperature in the BCS limit of a conventional two-component Fermi gas with $s$-wave scattering length $a_{s}<0$ [60]:
$T^{*}/T_{F}=\frac{8}{\pi}e^{\gamma-2}\begin{cases}\exp(\frac{\pi/c^{2}}{2k_{F}a_{13}}+\frac{2\pi^{2}}{mk_{F}}\frac{s^{2}}{c^{2}}I_{+}),&a_{23}=0,\\ \exp(\frac{\pi/s^{2}}{2k_{F}a_{23}}+\frac{2\pi^{2}}{mk_{F}}\frac{c^{2}}{s^{2}}I_{+}),&a_{13}=0.\end{cases}$ (48)
In Fig. 5, we show the critical temperature and total chemical potential for a set of negative scattering lengths in the BCS limit, at fixed $\Omega/E_{F}=4$. Here, we find a single solution corresponding to the ground state. We see that the critical temperature remains low but nonzero across the range of detunings, and that furthermore the numerical result for $T^{*}$ (solid black line) is very well approximated by Eq. (52) with $F=\frac{\pi}{2k_{F}a_{-}}$ (dashed red line). We also see that the chemical potential is close to that of free fermions.
IV.3 Thouless criterion in the limit of zero temperature
@@
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ +\sum_{{\bf k}}\left[\frac{\tanh(\beta\frac{\bar{\epsilon}_{{\bf k}}-2E_{F}}{2}\frac{m_{r}}{m_{3}})}{2(\bar{\epsilon}_{\bf k}-2E_{F})}-\frac{1}{2\bar{\epsilon}_{\bf k}}\right]$
@@
IV.4 Thouless criterion at finite temperature

### v2

@@ Abstract
We explore the paired superfluid phases of a Fermi gas in the presence of a continuous Rabi drive. We focus on the case where two components are Rabi coupled strongly enough that they form hybrid superpositions, which in turn interact with an uncoupled third component. Using a generalized Bardeen-Cooper-Schrieffer (BCS) ansatz, we show that there are two coupled superfluid order parameters, and we obtain the associated free energy and quasiparticle excitation spectrum. We find that we can drive BCS-BCS, BCS-Bose-Einstein condensate (BEC) and BEC-BEC crossovers purely by varying the detuning of the Rabi drive from the bare transition, with the precise crossover depending on the sign of the underlying interactions between the coupled and uncoupled components. We furthermore identify an exotic excited branch which features both normal to BCS superfluid transitions, as well as a BCS-BEC-BCS crossover. Introducing a generalized Thouless criterion, we show that this behavior is reflected in the critical temperature for superfluidity. Our Rabi-coupled scenario also possesses additional thermodynamic properties related to the pseudospin of the coupled components, which provide novel signatures of the state of the many-body system. The Rabi-driven Fermi gas thus emerges as a unique platform for engineering and probing a rich array of multi-band superfluid phases.
@@ I Introduction
Motivated by these developments, here we investigate the superfluid properties of a strongly interacting, strongly Rabi-coupled Fermi gas. The central idea is illustrated in Fig. 1. A continuous rf field creates a superposition (purple) of two atomic states ($|1\rangle$=red and $|2\rangle$=blue), where each state interacts differently with a third atomic species ($|3\rangle$=gray). Taking the total number of atoms in the Rabi-driven components equal to the third component, i.e., $N_{1}+N_{2}=N_{3}$, we find that there is the prospect of driving a rich array of different superfluid crossovers—BCS-BCS, BCS-BEC, and BEC-BEC—purely by varying the detuning of the rf field from the bare transition. While our proposed setup naturally involves three states, we emphasize that for sufficiently large Rabi coupling, where its strength exceeds the Fermi energy, the system effectively becomes a two-component Fermi gas where the upper Rabi-dressed state is minimally occupied. In this sense, our focus is distinct from previous theory works on three-component Fermi gases, such as those exploring analogs of color superfluidity [33, 34, 35, 36, 37, 38, 39, 40].
@@
[…]
@@ II Model
Here, $\hat{f}_{{\bf k}j}$ ($\hat{f}_{{\bf k}j}^{\dagger}$) are annihilation (creation) operators for fermions of species $j=\{1,2,3\}$. The states $|1\rangle$ and $|2\rangle$ are coupled via a Rabi drive which we treat within the rotating wave approximation. The drive has strength $\Omega$ and it is detuned by $\delta$ with respect to the bare $|1\rangle$-$|2\rangle$ transition. We therefore take the dispersions for each species at momentum $\bf k$} to be $\epsilon_{{\bf k}1}\equiv\epsilon_{{\bf k}}=k^{2}/2m$, $\epsilon_{{\bf k}2}=\epsilon_{\bf k}+\delta$ and $\epsilon_{{\bf k}3}=k^{2}/2m_{3}$. We take a general mass ratio $m_{3}/m$ for species $|3\rangle$ compared with the coupled species $|1\rangle$ and $|2\rangle$, since many of our results for the paired superfluid phases only depend on the reduced mass $m_{r}=mm_{3}/(m+m_{3})$. They thus apply to general mass-imbalanced Fermi mixtures [42], as well as equal-mass 6Li or 40K gases where the three species would represent three distinct hyperfine states. For instance, the scenario where three hyperfine components of 6Li are simultaneously present has already been experimentally realized in Refs. [43, 44, 45, 46, 47]. Throughout, we work in units where the reduced Planck constant $\hbar$, the Boltzmann constant $k_{\rm B}$, and the volume are all set to 1.
[…]
@@ III BCS mean field theory
$[\hat{H},\hat{N}_{3}]=0,\penalty 10000\ \penalty 10000\ [\hat{H},\hat{N}_{1}+\hat{N}_{2}]=0.$ (9)
@@ III.1 Generalized BCS ansatz
[…]
${\color[rgb]{0,0,0}\begin{pmatrix}\hat{f}_{{\bf k}a}\\ \hat{f}_{{\bf k}s}\end{pmatrix}=\begin{pmatrix}c_{\bf k}&&-s_{\bf k}\\ s_{\bf k}^{*}&&c_{\bf k}^{*}\end{pmatrix}\begin{pmatrix}\hat{f}_{{\bf k}1}\\ \hat{f}_{{\bf k}2}\end{pmatrix},}$ (11)
with $|c_{\bf k}|^{2}+|s_{\bf k}|^{2}=1$. Note that, unlike the exact coefficients $c$ and $s$ in Eq. (5), $c_{\bf k}$, $s_{\bf k}$ are variational coefficients that define the distribution between the two pairing states and thus remain to be determined [^1] In general, once the system is interacting we have $\hat{f}_{{\bf k}a/s}\neq\hat{f}_{{\bf k}\mp}$. Here, we use the indices $a,s$ by analogy with antisymmetric and symmetric superpositions..
@@ III.2 Gap equations
$0$ $=\zeta_{\bf k}u_{\bf k}v_{\bf k}+(u_{\bf k}^{2}-v_{\bf k}^{2})\left(\Delta_{1}c_{\bf k}-\Delta_{2}s_{\bf k}\right),$ (14a)
$0$ $=\left[2\delta c_{\bf k}s_{\bf k}-\Omega(c_{\bf k}^{2}-s_{\bf k}^{2})\right]v_{\bf k}^{2}-2u_{\bf k}v_{\bf k}\left(\Delta_{1}s_{\bf k}+\Delta_{2}c_{\bf k}\right).$ (14b)
Here, we have introduced the two real order parameters $\Delta_{1}$ and $\Delta_{2}$ defined by
@@
Note that, had we considered arbitrary phases of the Rabi drive and variational parameters, the phase difference between $\Delta_{1}$ and $\Delta_{2}$ would be locked to that of the Rabi drive, up to multiples of $\pi$. This would not change any of our conclusions below, and hence we keep all of these parameters real.
[…]
$v_{\bf k}^{2}=\frac{1}{2}\left(1-\frac{\zeta_{\bf k}}{2E_{\bf k}}\right),\penalty 10000\ \penalty 10000\ u_{\bf k}v_{\bf k}=-\frac{\Delta_{1}c_{\bf k}-\Delta_{2}s_{\bf k}}{2E_{\bf k}},$ (16)
@@
$\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \times\frac{\left(t_{\bf k}^{2}+2t_{\bf k}\frac{\delta}{\Omega}-1\right)}{4\left(t_{\bf k}+\frac{\Delta_{2}}{\Delta_{1}}\right)^{2}\left(1-t_{\bf k}\frac{\Delta_{2}}{\Delta_{1}}\right)}=\frac{\Delta_{1}^{2}}{\Omega^{2}},$ (22)
@@
In our model where we take $\Omega>0$, the solution with $\text{sgn}(\Delta_{2}/\Delta_{1})=-1$ corresponds to the stable ground state, while the one with $\text{sgn}(\Delta_{2}/\Delta_{1})=+1$ corresponds to an excited state, which appears as a saddle point of the free energy. This is illustrated in Fig. 2, which shows colormaps of $\mathcal{F_{\text{BCS}}}(\Delta_{1},\Delta_{2})$. Indeed, we can observe in panel (a) the existence of 4 stationary points corresponding to a pair of minima with $\text{sgn}(\Delta_{2}/\Delta_{1})=-1$, and a pair of saddle points appearing in the orthogonal direction in the $(\Delta_{1},\Delta_{2})$ plane. We see that the excited branch is topologically distinct from the ground state, in the sense that one of the order parameters must go through zero when evolving from one state to another. Furthermore, as is evident in Fig. 2, the free energy is very flat in the region around the saddle point. This suggests the possibility of tuning into a many-body excited state that is long lived compared to the typical timescales of the system.
@@
We also observe that the two types of BCS solutions are connected to the two-body bound states in the limit of vanishing density (see Appendix A). However, we stress that while it has been shown that the existence of two bound states requires both scattering lengths to be positive [41], we find that this is not a necessary condition for two solutions of the order parameters to exist in the many-body regime. In Section IV, we discuss how one can use a generalized Thouless criterion to investigate the existence of the excited state in the many-body regime.
@@ III.3 Quasiparticle excitations
[…]
@@ IV Thouless criterion
$\penalty 10000\ \penalty 10000\ -\left(\Delta_{1}\quad\Delta_{2}\right)\mathbf{T}^{\text{mb}}(0)^{-1}\begin{pmatrix}\Delta_{1}\\ \Delta_{2}\end{pmatrix}+\mathcal{O}(\Delta_{i}^{2}\Delta_{j}^{2}).$ (36)
@@
This corresponds to a generalized version of the Thouless criterion [58]. We note that a Thouless criterion given by the determinant of a two-by-two matrix also occurs in the model of two-band superfluids used in Refs. [59, 60]. However, this model differs from the one we consider since it describes a four-component Fermi system, while we consider a system with three components. Furthermore, in the models of two-band superfluids [52, 53, 59, 60], the interband coupling has a two-body origin and couples pairs from different bands, whereas here the Rabi coupling is a single-particle effect.
Solving Eq. (38) together with the number equations in Eq. (35) allows us to determine the critical temperature at fixed density. Specifically, renormalizing the diagonal elements of the inverse many-body $T$ matrix in Eq. (37) in the same manner as the gap equation in Eq. (19), we find that the Thouless criterion takes the form
@@
$\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ +\left(s^{2}\frac{m_{r}}{2\pi a_{13}}+c^{2}\frac{m_{r}}{2\pi a_{23}}\right)I_{-}+I_{+}I_{-}=0,$ (39)
@@ IV.1 Thouless criterion in the limit of zero temperature
$\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ +\sum_{{\bf k}}\left[\frac{\tanh(\beta\frac{\bar{\epsilon}_{{\bf k}}-2E_{F}}{2}\frac{m_{r}}{m_{3}})}{2(\bar{\epsilon}_{\bf k}-2E_{F})}-\frac{1}{2\bar{\epsilon}_{\bf k}}\right]$
@@
IV.2 Thouless criterion at finite temperature